Please someone help I’m stick

Two women have decided to start a cottage industry that will be run out of their homes. They are going to make two types of specialty shirts, rhinestone and embroidery. The total gross profit is given by

0 0
Answer #1

oL and d= 20-2 16-2n = 20-24 2 2. 2L7 Tw.ghineNone xhi t a-d 4 emb.niard